Given 35 degrees (upper) and 7 degrees (lower)
with the following equations to substitute,
1. x = 14-d/tan(7)
2. x = d/tan(35)
along with x = 14/tan(7)+tan(35)
i am getting 24.294367 (17.01109891/tan(35)) and -24.52343265(-3.01109891/tan(7)). i got the -3.01109891 by subtracting 14 with 17.01109891. They should both be the same answer. I am sure about the first answer but not the second, so i would really appreciate your help.
Cheers

Answer by mananth(16946)


If my sketch is right

x/y = tan 7
y= x/tan 7

(x+14)/tan 35 = y

x/tan 7 = (x+14)/tan35
tan 35*x = tan7*x +tan 7*14
x(tan 35-tan7) = tan 7 * 14
x = 1.71/ (tan 35-tan 7)
x= 1.71/0.577
x= 2.96
Height of hill =( 14+2.96-1.73)
=15.23
y= 2.96/tan 7
y = 24.10 horizontal distance

Check
y = 16.96/tan 35
=24.20